OpenStudy (anonymous):

Is cos(1/(x+1)) ever undefined?

OpenStudy (anonymous):

When x = -1

OpenStudy (anonymous):

But when I input that into my calculator it says cos(0)=1...is that still undefined?

OpenStudy (anonymous):

well, 1/0 is not equal to 0 actually so you can't say cos (1 / 0) = cos(0)

OpenStudy (anonymous):

1/0 = undefined

OpenStudy (anonymous):

you, are a genius.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

haha, not a problem my friend...have fun!